the circle diameter is 26cm there is a rectangle in the circle and the short side of the rectangle is 10cm how do tou find the long side of the rectangle???
Answer by dfrazzetto(283)


The diagonal of the rectangle HAS to be the diameter of the circle, in this case 26cm, so:
It is a right triangle:
a=10cm
b=?
c=hypotenuse= 26
*using Pythagorean Theorem*
a^2+b^2=c^2
10^2+b^2=26^2
100 = b^2 = 676
b^2 = 576
b= 24 cm